OpenStudy (anonymous):

The area of trapezoid can be found using the formula 1/2h(b+B) if the area of trapezoid is 100 sq in. With a height of 10 and base of 12 determined. The length of the base

OpenStudy (anonymous):

@wateriyenes

OpenStudy (allank):

Hint: 100=0.5*10*(12+B)

OpenStudy (anonymous):

100=60b?

OpenStudy (anonymous):

@allank

OpenStudy (anonymous):

how'd u get that 60 on the right side?

OpenStudy (anonymous):

10x12+2?

OpenStudy (anonymous):

not quite... according to the hint given by allank, you need to multiply the 0.5 * 10 first, then distribute: \(\large 100=\color {red}{(\frac{1}{2})(10)}(12+B) \) \(\large 100=(5)(12+B) \)

OpenStudy (anonymous):

can you figure out what the next line would be from here? hint: distribute the right side...

OpenStudy (anonymous):

100=60+b

OpenStudy (anonymous):

not quite.... try again... just working on the right side, 5(12+B) =

OpenStudy (anonymous):

Is it 40?

OpenStudy (anonymous):

\(\large (5)(12+B)=5\cdot 12 + 5 \cdot B \rightarrow 60+5B \) that is what you should have on the right side.... now replace that right side with the one i have and solve....

OpenStudy (anonymous):

solve: \(\large 100=60+5B \)

OpenStudy (anonymous):

40/5?

OpenStudy (anonymous):

simplify....

OpenStudy (anonymous):

B = 40/5 = ???

OpenStudy (anonymous):

The answer is 8

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw..